OpenStudy (anonymous):

convert the angle measures to degrees 1. pi/6 2. 3pi/5

OpenStudy (anonymous):

The conversion between degrees and radians (and vice-versa) is\[180 \text{ deg}=\pi\text{ rad}\] What happens you divide both sides by 6, for the first question?

OpenStudy (phi):

do the opposite of the other problems multiply by \[\frac{180 º}{\pi \text{ radians}} \]

OpenStudy (phi):

notice you can cancel the pi right away \[ \frac{\cancel{\pi}}{6} \cdot \frac{180}{\cancel{\pi}} = \frac{180}{6}\]

OpenStudy (phi):

the way to remember what to multiply is write down the units: \[ \frac{\pi \text{ radians}} {6} \] if you want degrees, they you want to multiply by degrees over radians , so you know you want to use \[ \frac{\pi \text{ radians}} {6} \cdot \frac{180 \text{ degrees}}{\pi \text{ radians}} \] because the radians will cancel \[ \frac{\cancel{\pi \text{ radians}}} {6} \cdot \frac{180 \text{ degrees}}{\cancel{\pi \text{ radians}}} \]

OpenStudy (anonymous):

@phi is that for pi/6?

OpenStudy (anonymous):

π/6 x 180/ π = 180/6 is that the answer for pi/6 or does it continue into the next post you wrote?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

I only did pi/6. of course 180/6 simplifies to 30 degrees, right ?

OpenStudy (anonymous):

yes, so in that case is 30 the final answer? sorry i have to post this and then go over the answer

OpenStudy (phi):

yes. People sometimes memorize pi/6 rads is 30º

OpenStudy (anonymous):

thanks so much :) so pi/6=30 degrees?

OpenStudy (anonymous):

really appreciate it youre great at this! :)

OpenStudy (phi):

did you get the answer to the 2nd problem ?

OpenStudy (anonymous):

ive gotta go now but ill type it here later and maybe u can check it if u get a chance :) thanks once again!!